THE ENERGY LEVELS OF LANTHANIDE IONS, FREE AND DOPED INTO CALCIUM FLUORIDE

\underline{SCOTT R. BROZELL} and RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Lanthanide ions, both free and in CaF_{2} crystals, have been studied extensively via optical spectroscopy and electron paramagnetic resonance. Ab initio spin-orbit configuration interaction calculations are performed on some of these systems. The primary properties of interest are the energy levels and magnetic moments. The CaF_{2} host is modeled with a large finite cluster of ions which approximate the Madelung potential of the crystal lattice. Lanthanide ions are modeled with relativistic effective core potentials and Gaussian cc-pVDZ basis sets. The results for the lanthanide ions are compared and contrasted with similar studies on actinide ions.